Grade 12th passMechanics

a projectile is projected with speed u at an angle of 60 degree with horizental from the foot an inclined plane if the projectile hites the inclined plane horizontally the range on inclined plane wiil be

naveen
10 Years agoGrade 12th pass

1 Answer

Aakanksha Talreja
10 Years ago
projectile hits the plane horizontally \Rightarrow v_{y}=0 \thereforeusin(\alpha - \beta) – gsin\betat = 0
t= 2u\sin (\alpha - \beta ) / g\cos \beta
u\sin \left ( \alpha -\beta \right )= g\sin \beta 2u\sin \left ( \alpha -\beta \right ) / g\cos \beta \Rightarrow \tan \beta = 1/2
range R= 2u^{2}\sin \left ( \alpha -\beta \right )\cos \alpha /g\cos ^{2}\beta
R=u^{2} (2\sqrt{15} - \sqrt{5}) / 80